OpenStudy (anonymous):

The length of a rectangle is 7 mm longer than its width. Its perimeter is more than 62 mm. Let w equal the width of the rectangle. a.Write an expression for the length in terms of the width b.Use these expressions to write an inequality based on the given information. c.Solve the inequality, clearly indicating the width of the rectangle.

jhonyy9 (jhonyy9):

l=7+w P>62 P=2l+2w=2(7+w)+2w=14+2w+2w=14+4w 14+4w>62 4w>48 w>48/4 w>12